OpenStudy (anonymous):

Tracey uses one-fourth of a cup of brown sugar in each batch of cookies. If she makes thirteen batches of cookies, how much brown sugar will she use? three and three-fourths three and one-fourth three and one-third three and one-half

OpenStudy (anonymous):

as 13 is a whole number multiply the top number of the fraction by 13 = 13/4 then simplify by dividing the top number by the bottom number to give whole numbers and the remainder is still a fraction = 3 (as 3 x 4 = 12) and 1/4 left over = 3 and 1/4